CH/Agen/95/01

Agenda

Charissa Group Meeting

To be held on: Wednesday, 10 May 1995, 10.30 am

To be held at:  York, Room PT103

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

Agenda

Begin Charissa Meeting, 10.30 a.m., Room PT103 View MEGHA system 12.20 p.m. Buffet Lunch 1.00 p.m.

Conclude Charissa Meeting, 2.00–3.30 p.m., Room PT103

1. Apologies
2. Minutes of 26 September 1993 (meet9505)
3. Matters Arising / Actions
4. Update on EPSRC matters

Discussion on review and ANU review

5. Progress with MEGHA development

i Hardware (detectors, chamber, gas supply)

ii Electronics

iii Data Acquisition

6. Installation of MEGHA at the ANU

i Problems with beamline shipment

ii Timescale for installation

7. Next ANU running

i Time

ii Experiments

iii Changes to hardware, electronics, data acquisition, software

8. Status of Strasbourg plans
9. Reports of other running

Orsay, Ganil, Argonne

10. Publications and Talks
11. Any other business
12. Date of next meeting

W.N. Catford (Secretary) 4 May 1995